# Graph Neural Backdoor: Fundamentals, Methodologies, Applications, and Future Directions

### v1

@@ 3. Current Methodologies
Based on this classification, we will first elaborate on the existing GNN backdoor attack research. After that, we will show the corresponding defensive strategy studies. The mind map for the current studies is illustrated in Fig
Figure 4. The mind map for current graph neural backdoor studies.
3.1. General Framework of GNN backdoor
For GNN model, adversaries typically resort to data-poisoning on the training data to implant backdoors. This manipulation will cause the well-trained GNN to produce adversary-designated outputs when input trigger-embedded graph data during deployment. Nevertheless, for clean inputs (unpoisonous or unattacked), the model will maintain its accuracy. The general framework of GNN backdoor is illustrated in Fig. 5.
@@ 3.1. General Framework of GNN backdoor
(7) $\mathcal{D}_{t}\sim\mathcal{D},\quad|\mathcal{D}_{t}|=\gamma|\mathcal{D}|,$
(8) $\mathcal{D}_{t}=\{(x_{i}=x_{i}+\Delta,y_{t})\,|\,(x_{i},y_{i})\in\mathcal{D}_{t}\}.$
@@
(9) $\text{GNN}_{\theta}=\underset{(x_{i},y_{i})\in\mathcal{D}}{\text{argmin}_{\theta}}\mathcal{L}(\text{GNN}(x_{i};\theta),y_{i}),$
@@
(10) $\text{GNN}(x_{i}+\Delta)=y_{t},\quad\text{GNN}(x_{i})=y_{i}.$
Figure 5. Illustration of the general process of GNN backdoor, which is achieved by data-poisoning. Specifically, the adversary opts for a subset of samples from the training data and inserts designated subgraphs as triggers into them, subsequently modifying the ground truths of these data to the target class. As a consequence, this causes the trained model to predict the target class for input samples embedded with these triggers.
3.2. Attack Methodologies
Based on the classification of backdoor attacks studies on GNN outlined at the beginning of this section, we will elaborate on them in two parts: (i) adaptability-expanding research; (ii) effectiveness-improving research. The comprehensive summary of backdoor attacks is shown in Tab. 3.
Table 3. Summary of Graph Neural Network Backdoors
Graph Node Edge Samples Model Loss
Subgraph Backdoor (2021) (Zhang et al., 2021) $\bullet$ - - $\bullet$ - - General GNN Seminal paper
GTA (2021) (Xi et al., 2021) $\bullet$ $\bullet$ - $\bullet$ - $\bullet$ General GNN Seminal paper
Random Trigger Backdoor (2021) (Sheng et al., 2021) $\bullet$ - - $\bullet$ - - General GNN Optimize attack node selection
CBA/DBA (2022) (Xu et al., 2022, 2023a) $\bullet$ - - $\bullet$ $\bullet$ - Federated GNN Extend backdoor to FL
Bkd-FedGNN (2023) (Liu et al., 2023) $\bullet$ $\bullet$ - $\bullet$ - $\bullet$ Federated GNN Extend FL backdoor to node-level task
GCBA (2023) (Zhang et al., 2023a) - $\bullet$ - $\bullet$ - - Contrastive learning GNN Extend backdoor to contrastive learning
CrossBA (2024) (Lyu et al., 2024) - $\bullet$ - $\bullet$ - $\bullet$ Graph prompt learning Extend backdoor to graph prompt learning
PoisonedGNN (2023) (Alrahis et al., 2023) $\bullet$ - - $\bullet$ $\bullet$ $\bullet$ Hardware GNN system First physical GNN system backdoor
Explainability Backdoor (2020) (Xu et al., 2021) $\bullet$ $\bullet$ - $\bullet$ $\bullet$ - General GNN Employ explainability to design trigger
Motif Backdoor (2023) (Zheng et al., 2023) $\bullet$ - - $\bullet$ - - General GNN Explain GNN Backdoor Effectiveness
General Backdoor (2022) (Chen et al., 2022) - $\bullet$ - $\bullet$ - $\bullet$ General GNN Employ explainability to improve stealthiness
Explanatory Subgraph Attack (2024) (Wang et al., 2024c) $\bullet$ - - $\bullet$ $\bullet$ $\bullet$ General GNN Keep explanatory information for poisoned graphs
UGBA (2023) (Dai et al., 2023a) - $\bullet$ - $\bullet$ $\bullet$ $\bullet$ General GNN Optimize to improve stealthness
CBAG (2024) (Dai and Sun, 2024b) - $\bullet$ - $\bullet$ - - General GNN Propose clean-graph backdoor
Poster (2023) (Xu and Picek, 2022) $\bullet$ - - $\bullet$ - - General GNN Propose clean-label GNN backdoor
CGBA (2023) (Xing et al., 2023) - $\bullet$ - $\bullet$ - - General GNN Reduce poisonous perturbation
PerCBA (2023) (Yang et al., 2023a) - $\bullet$ - $\bullet$ - - General GNN Extend clean-label backdoor to node-level task
TRAP (2022) (Yang et al., 2022) $\bullet$ - - $\bullet$ $\bullet$ $\bullet$ General GNN Employ data transferability to design backdoor
Graph Spectrum Backdoor (2023) (Zhao et al., 2023) $\bullet$ - - $\bullet$ - - General GNN Employ data spectrum to design backdoor
NFTA (2023) (Chen et al., 2023) $\bullet$ - - $\bullet$ - - General GNN Improve performance by smoothing out triggers
Link-Backdoor (2023) (Zheng et al., 2024) - - $\bullet$ $\bullet$ - $\bullet$ General GNN Extend backdoor to edge-level task
Link Prediction Attack (2024) (Dai and Sun, 2024a) - - $\bullet$ $\bullet$ - - General GNN Improve edge backdoor triggers
MLGB (2024) (Wang et al., 2024a) - $\bullet$ - $\bullet$ $\bullet$ $\bullet$ General GNN Introduce N-to-N multi-target backdoor attack
Poster (2023) (Xu and Picek, 2023) - $\bullet$ - $\bullet$ - - General GNN Introduce 1-to-N&N-to-1 multi-target backdoor
3.2.1. Adaptability-expanding Research
@@
Pioneering Works. Zhang et al. initially investigated the feasibility of embedding backdoors within GNNs (Zhang et al., 2021). This was achieved through using subgraphs as triggers, which were strategically inserted into the graph structures of a randomly selected subset of training samples, and simultaneously, altering the source labels of these samples to the target class (e.g., data-poisoning). The poisoned training set (including poisoned subset and clean subset) was then utilized for GNN learning, and this led to the trained model becoming infected (backdoored), resultantly causing misclassification of trigger-embedded graph inputs into the adversary’s predetermined target class. Notably, the model still maintained its accuracy in classifying regular graph samples amidst this backdoor affection.
This study further explores the impact of different categories of subgraph triggers on the effectiveness of the backdoor. It introduces two types of trigger designs: fixed-pattern subgraphs and generated-pattern subgraphs. While the fixed-pattern subgraph mode (e.g., complete subgraph) is relatively singular and easier to detect and filter, a generation algorithm (e.g., preferential attachment model) is employed to create diverse subgraphs, enhancing the stealthiness of the attack. Moreover, (Sheng et al., 2021) presented a similar attack scheme and optimizes the selection of target subgraph for trigger insertion by degree centrality and closeness centrality.
Another pioneering research in GNN backdoor is Graph Trojaning Attack (GTA) (Xi et al., 2021) [^1] Both papers of GTA and Subgraph Backdoor in (Zhang et al., 2021) claim to be the first exploitation research in GNN backdoor, but in terms of the publication order, Subgraph Backdoor should be earlier.. It also leverages data-poisoning to infect the training dataset to embed backdoor. However, in contrast to Subgraph Backdoor, it differs in two key aspects: trigger design and model training. Firstly, it doesn’t employ fixed models of trigger generation for backdoor implantation. Instead, it utilizes a dynamically-trained backdoor trigger generator to craft subgraph trigger tailored to individual input graph structure. Secondly, it introduces a bi-level optimization training paradigm. This involves concurrently training the model and the trigger generator, which not only confirms the normal training of the model but also enforces the generator to create corresponding trigger subgraphs based on the features of input graph structures, rather than relying on fixed subgraph-generation model or algorithm.
@@
Xu et al. first explored the possibility of implanting backdoors into this paradigm and presented two types of data-poisoning-based strategies: Distributed Backdoor Attack (DBA) and Centralized Backdoor Attack (CBA) (Xu et al., 2022, 2023a).
  1. (1)
@@
Liu et al. extended the backdoor to node classification scenarios via a stated local loss function (Liu et al., 2023). Additionally, they discussed various factors influencing the backdoor performance in federated graph learning (e.g., data distribution, adversary client number, trigger size, location, poisoning rate, etc.).
@@
To implant backdoor into the GCL framework, Zhang et al. first looked into the viability and proposed three methods according to the attack scenarios: (i) data-poisoning attack; (ii) set-crafting attack; (iii) natural backdoor attack (Zhang et al., 2023a).
  1. (1)
    Data-poisoning Attack. This attack confines adversaries to control the training data collection process of GCL. One core concept of GCL involves training to maximize the similarity between embeddings of the same graph node across different mapped views. Hence, by selecting a subset of nodes from the target class in the graph and connecting them to nodes with triggers, certain connections may disappear in some views after mapping through the encoder (nodes won’t aggregate trigger node information), while others will remain (nodes will aggregate trigger node information). Consequently, when calculating the loss, the encoder training aims will maximize the similarity between embeddings with trigger information and those without, hence achieving the effect of a backdoor.
  2. (2)
@@
Graph Prompt Backdoor. Applying pre-trained GNNs to downstream tasks that differ from the pre-training stage poses challenges related to generalization and knowledge transfer. Graph Prompt Learning (GPL) emerges as a solution by training the GNN encoder on unlabeled data and customizing prompts for downstream tasks to effectively bridge the gap between pre-training and downstream tasks and promote knowledge transfer.
Research (Lyu et al., 2024) proposes a CrossBA attack for GPL. It aims to introduce embedding collisions between backdoored graphs and the trigger graph within the backdoored GNN encoder while ensuring that the backdoored graph maintains a distinct embedding compared to its clean counterpart. This backdoor poisoning effect results in the backdoored graph being associated with a target embedding that is distant from the clean input, leading to misclassification in downstream models. To achieve this, the model is trained using four subloss functions: (1) main learning loss, which defines the learning objective of the main task; (2) backdoor learning loss, which defines the backdoor task objective, utilizing the embedding of the trigger graph as the target, and induces the backdoor mapping into the GNN encoder by colliding the embeddings of backdoored graphs with that of the trigger graph; (3) embedding alignment loss, which is designed to ensure that backdoors do not impair the GNN encoder’s ability to generate discriminative embeddings for clean data; and (4) node feature affinity loss, which is designed to enhance stealthiness to evade anomaly checks.
@@
Research (Alrahis et al., 2023) initially investigated the GNN backdoor in integrated circuits, utilizing subcircuits as triggers to infect the circuit dataset used for training. The embedded subcircuits do not affect the performance of the original circuit and can pass functional convergence tests. The backdoored GNN physical system classifies the circuit with triggers as harmful. Experimental results based on Hardware Trojan Detection Systems and Intellectual Property Piracy Detection Systems demonstrate that the proposed backdoor scheme exhibits excellent performance.
3.2.2. Effectiveness-improving Research.
@@
Xu et al. proposed an Explainability-based Backdoor Attack, exploiting node features with either the strongest or weakest characteristics to design subgraph triggers for poisoning the training data and embedding backdoors (Xu et al., 2021). Executing attacks using data with prominent features makes the GNN model efficiently learn the corresponding trigger features, and facilitates more effective backdoor implantation. Conversely, launching backdoor attacks using data with weaker features minimally affects the feature learning process for normal samples, rendering the attack more covert and less prone to detection.
@@
Zheng et al. researched the distribution patterns in statistical terms between trigger subgraphs and normal subgraphs (Zheng et al., 2023). Their findings indicated that utilizing subgraphs divergent from the distribution of normal samples as triggers resulted in a notable enhancement in attack performance. Moreover, based on this phenomenon, they proposed a motif-based backdoor poisoning technique. This method employs available data to select significantly skewed subgraph distributions as motif triggers and applies node importance and subscore to determine the optimal trigger injecting position with the highest values. Then, the adversary can poison the training data with the designed trigger and backdoor the associated GNN models via learning.
Chen et al. introduced a general GNN backdoor tailored for node-level tasks (Chen et al., 2022). Employing a trigger generator, the attack generates constrained node features specifically designed for poisoned targets, with both the GNN model and the generator being jointly trained via the poisonous dataset, which resultantly creates a backdoored GNN. Furthermore, to reduce the impact of other nodes, it selectively eliminates critical edges of targets, identified through explainability analysis. Additionally, the trigger feature dimensions crucial for targets are prioritized.
Research (Wang et al., 2024c) proposed an explanatory subgraph-based poisoning attack. The subgraph incorporates graph structure and node feature information relevant to the target label, and this strengthens the association between the subgraph and the target while minimally altering the original data’s structure, and preserving its statistical properties. Also, the adversary will modify the explanation information of the graph, which makes the change of the graph more covert and reasonable.
@@
Dai et al. developed imperceptible trojan backdoor attacks under limited attack budgets (Dai et al., 2023a). Utilizing a trigger generator, the generation of infective samples ensures a sufficiently high similarity with poisoned samples from the target class data. Additionally, to make the model maintain the training accuracy on normal samples, a bi-optimization loss is designed to train both trigger generator and GNN. Furthermore, for the selection of the target for the attack, the method leverages the obtained node embeddings for clustering and chooses the most representative nodes as targets to assure backdoor efficiency.
To achieve attack concealment in graphs, study (Dai and Sun, 2024b) developed a clean-graph backdoor attack. This attack refrains from modifying the graph’s features or structure and does not introduce malicious nodes and edges prior to training. Instead, adversaries solely alter the labels of victim nodes possessing adversary-desired distinctive features. By doing so, they ensure that the GNN training process leads the model to produce target outcomes when presented with input data exhibiting these features.
@@
Xu et al. initially delved into the feasibility of implementing clean-label backdoor attacks in GNNs (Xu and Picek, 2022). It samples victim targets within the training dataset pertaining to the target class, inserting triggers within them, and then training the GNN model using both poisoned and regular data. Since the infected data belongs to the target class, it suffices to embed triggers in these samples to enable the GNN to learn trigger features without label revision. During the testing phase, the GNN will classify trigger-contained inputs as belonging to the target class. To enhance this method, Xing et al. select robust features of nodes with larger degrees as triggers (Xing et al., 2023). These features are typically strongly associated with their respective labels. Utilizing them as triggers facilitates the model in establishing a strong correlation between the trigger and the target label.
Yang et al. evaluated the potential of clean-label attacks for node classification (Yang et al., 2023a). To realize the backdoor effect through poisoning without label modification, they embedded triggers within the target node data and introduced perturbations therein. This action causes the decision boundary of the target category to lean toward infected data during training, and leads the model to predict attacked data as the target class during the testing phase. Simultaneously, to enhance the normal performance robustness of GNN, they also proposed a contrastive poisoning strategy to slightly poison normal samples.
@@
The transferable GNN backdoor attack is proposed in (Yang et al., 2022), utilizing available real-world data to create triggers for poisoning and implanting a backdoor in a shadow GNN model via training. The mode of subgraph triggers is optimized during training using loss function derivatives to maximize the attack success rate, and it could be applied in poisoning various downstream task GNN models (i.e., transferability). Additionally, due to the harsh computing of optimizing graph structure derivatives, a solution strategy for graph structure perturbations was also presented in the study.
@@
Inspired by frequency domain graph theory and frequency-based backdoor attack methods, Zhao et al. proposed an effective spectral graph backdoor attack method (Zhao et al., 2023). It injects subtle trigger signals into the crucial frequency bands of important node features to poison training samples and implant backdoors within victim GNN. For the trigger, specifically, it first extracts the ego graph constructed from the attributes of the neighboring nodes of the target node and performs spectrum decomposition on this ego graph. Then, the adversaries inject a small trigger signal into the spectrum graph after the spectrum decomposition. The injection is done at the important frequency bands of the important features, to guarantee an efficient backdoor effect.
Graph Node Feature Backdoor. A novel backdoor approach is proposed based on graph node features, aiming to infect nodes and optimize the smoothness of graph features to generate an infected graph for backdooring GNNs (Chen et al., 2023). Concretely, the attack employs a mask to conceal node feature data and derive features. Following that, a graph feature optimization function is utilized to modify the graph structure, resulting in the smoothest infected graph containing the trigger. The features of the graph are smoother, minimizing the impact on feature learning from the original data. The backdoor keeps better preservation of the model’s accuracy on normal samples while ensuring the robustness of the backdoor performance.
Graph Link Backdoor. Link prediction is an essential task in GNN, primarily targeting at predicting links (edges or connections) that are not yet present but may appear in the future within a graph. Zheng et al. initially analyzed the feasibility of implementing backdoors in such tasks, exploiting poisoning to backdoor GNN (Zheng et al., 2024). It first designates a specified subgraph structure as the trigger, then optimizes it using gradient information from the prediction model, and finally poisons the training data with the optimized trigger to backdoor the victim GNN. Dai et al. further improved the trigger by leveraging nodes as the trigger. During training, edges between nodes connected to this trigger are misclassified into an adversary-specified target (Dai and Sun, 2024a).
@@
Study (Wang et al., 2024a) reviewed the possibility of implementing such attacks within graph learning framework. First, explainability techniques are applied to determine multi-trigger locations and identify poisoned graph nodes according to these locations. Next, a one-hot encoding mechanism is used to encode the poisoned nodes and multiple target labels. Finally, one carefully designed loss function, coupled with a two-tier-optimization mechanism, is leveraged to train the multi-trigger generation model and victim GNN, resulting in the GNN backdoored that responds to multi-triggers after learning.
Xu et al. implemented multi-target backdoors through multi-target data-poisoning (Xu and Picek, 2023) [^2] The methods proposed in (Xu and Picek, 2023) and (Xu and Picek, 2022) are both named “Poster”.. The attack primarily centers on two backdoor strategies: One-to-N, in which distinct numerical values of several feature dimensions regulate various target outcomes, and N-to-One, whereby when all the triggers are satisfied, the backdoor could be activated.
Table 4. Summary of Graph Neural Backdoor Defense Strategies
@@
BloGBaD (2024) (Yang et al., 2024) $\bullet$ - - $\bullet$ - $\bullet$ $\diamond$ Backdoor detection $\diamond$ Sample filtration $\diamond$ Backdoor mitigation Employ sample distributions to identify infected graphs
Explainability Defense (2022) (Jiang and Li, 2022) $\bullet$ - - $\bullet$ $\bullet$ - $\diamond$ Backdoor detection $\diamond$ Sample filtration Leverage explainability tools to detect poisonous graphs
Securing GNN (2024) (Downer et al., 2024) $\bullet$ - - $\bullet$ $\bullet$ - $\diamond$ Backdoor detection Use graph features to detect sample anomalies
3.3. Defense Methodologies
In response to the threat of GNN backdoor attacks, researchers have begun focusing on developing effective defense strategies. The comprehensive summary of GNN backdoor defense studies is displayed in Tab. 4.

### v2

@@ III Backdoor Attack & Defense
Based on this classification, we will first elaborate on the existing GNN backdoor attack research. After that, we will show the corresponding defensive strategy studies.
TABLE I: Summary of Graph Neural Network Backdoors
Graph Node Edge Samples Model Loss
Subgraph Backdoor (2021) [15] $\bullet$ - - $\bullet$ - - General GNN Seminal paper
GTA (2021) [16] $\bullet$ $\bullet$ - $\bullet$ - $\bullet$ General GNN Seminal paper
Random Trigger Backdoor (2021) [19] $\bullet$ - - $\bullet$ - - General GNN Optimize attack node selection
CBA/DBA (2022) [20, 21] $\bullet$ - - $\bullet$ $\bullet$ - Federated GNN Extend backdoor to FL
Bkd-FedGNN (2023) [22] $\bullet$ $\bullet$ - $\bullet$ - $\bullet$ Federated GNN Extend FL backdoor to node-level task
GCBA (2023) [23] - $\bullet$ - $\bullet$ - - Contrastive learning GNN Extend backdoor to Contrastive learning
PoisonedGNN (2023) [24] $\bullet$ - - $\bullet$ $\bullet$ $\bullet$ Hardware GNN system First physical GNN system backdoor
Explainability Backdoor (2020) [25] $\bullet$ $\bullet$ - $\bullet$ $\bullet$ - General GNN Employ explainability to design trigger
Motif Backdoor (2023) [26] $\bullet$ - - $\bullet$ - - General GNN Explain GNN Backdoor Effectiveness
General Backdoor (2022) [27] - $\bullet$ - $\bullet$ - $\bullet$ General GNN Employ explainability to improve stealthiness
Explanatory Subgraph Attack (2024) [28] $\bullet$ - - $\bullet$ $\bullet$ $\bullet$ General GNN Keep explanatory information for poisoned graphs
UGBA (2023) [29] - $\bullet$ - $\bullet$ $\bullet$ $\bullet$ General GNN Optimize to improve stealthness
CBAG (2024) [30] - $\bullet$ - $\bullet$ - - General GNN Propose clean-graph backdoor
Clean-Label Backdoor (2023) [31] $\bullet$ - - $\bullet$ - - General GNN Propose clean-label GNN backdoor
CGBA (2023) [32] - $\bullet$ - $\bullet$ - - General GNN Reduce poisonous perturbation
PerCBA (2023) [33] - $\bullet$ - $\bullet$ - - General GNN Extend clean-label backdoor to node-level task
TRAP (2022) [34] $\bullet$ - - $\bullet$ $\bullet$ $\bullet$ General GNN Employ data transferability to design backdoor
Graph Spectrum Backdoor (2023) [35] $\bullet$ - - $\bullet$ - - General GNN Employ data spectrum to design backdoor
NFTA (2023) [36] $\bullet$ - - $\bullet$ - - General GNN Improve performance by smoothing out triggers
Link-Backdoor (2023) [37] - - $\bullet$ $\bullet$ - $\bullet$ General GNN Extend backdoor to edge-level task
Link Prediction Attack (2024) [38] - - $\bullet$ $\bullet$ - - General GNN Improve edge backdoor triggers
MLGB (2024) [39] - $\bullet$ - $\bullet$ $\bullet$ $\bullet$ General GNN Introduce N-to-N multi-target backdoor attack
Multi-Target Backdoor (2023) [40] - $\bullet$ - $\bullet$ - - General GNN Introduce 1-to-N&N-to-1 multi-target backdoor
III-A Framework of GNN backdoor
For GNN model, adversaries typically resort to data-poisoning on the training data to implant backdoors. This manipulation will cause the well-trained GNN to produce adversary-designated outputs when input trigger-embedded graph data during deployment. Nevertheless, for clean inputs (unpoisonous or unattacked), the model will maintain its accuracy.
@@ III-A Framework of GNN backdoor
$\mathcal{D}_{t}\sim\mathcal{D},\quad|\mathcal{D}_{t}|=\gamma|\mathcal{D}|,$ (5)
$\mathcal{D}_{t}=\{(x_{i}=x_{i}+\Delta,y_{t})\,|\,(x_{i},y_{i})\in\mathcal{D}_{t}\}.$ (6)
@@
$\text{GNN}_{\theta}=\underset{(x_{i},y_{i})\in\mathcal{D}}{\text{argmin}_{\theta}}\mathcal{L}(\text{GNN}(x_{i};\theta),y_{i}),$ (7)
@@
$\text{GNN}(x_{i}+\Delta)=y_{t},\quad\text{GNN}(x_{i})=y_{i}.$ (8)
III-B Attack Methodologies
Based on the classification of backdoor attacks studies on GNN outlined at the beginning of this section, we will elaborate on them in two parts: (i) adaptability-expanding research; (ii) effectiveness-improving research. The comprehensive summary of backdoor attacks is shown in Tab. I.
III-B1 Adaptability-Expanding Research
@@
Pioneering Works. Zhang et al. initially investigated the feasibility of embedding backdoors within GNNs [15]. This was achieved through using subgraphs as triggers, which were strategically inserted into the graph structures of a randomly selected subset of training samples, and simultaneously, altering the source labels of these samples to the target class (e.g., data-poisoning). The poisoned training set (including poisoned subset and clean subset) was then utilized for GNN learning, and this led to the trained model becoming infected (backdoored), resultantly causing misclassification of trigger-embedded graph inputs into the adversary’s predetermined target class. Notably, the model still maintained its accuracy in classifying regular graph samples amidst this backdoor affection.
This study further explores the impact of different categories of subgraph triggers on the effectiveness of the backdoor. It introduces two types of trigger designs: fixed-pattern subgraphs and generated-pattern subgraphs. While the fixed-pattern subgraph mode (e.g., complete subgraph) is relatively singular and easier to detect and filter, a generation algorithm (e.g., preferential attachment model) is employed to create diverse subgraphs, enhancing the stealthiness of the attack. Moreover, [19] presented a similar attack scheme and optimizes the selection of target subgraph for trigger insertion by degree centrality and closeness centrality.
Another pioneering research in GNN backdoor is Graph Trojaning Attack (GTA) [16]. It also leverages data-poisoning to infect the training dataset to embed backdoor. However, in contrast to Subgraph Backdoor, it differs in two key aspects: trigger design and model training. Firstly, it doesn’t employ fixed models of trigger generation for backdoor implantation. Instead, it utilizes a dynamically-trained backdoor trigger generator to craft subgraph trigger tailored to individual input graph structure. Secondly, it introduces a bi-level optimization training paradigm. This involves concurrently training the model and the trigger generator, which not only confirms the normal training of the model but also enforces the generator to create corresponding trigger subgraphs based on the features of input graph structures, rather than relying on fixed subgraph-generation model or algorithm.
@@
Xu et al. first explored the possibility of implanting backdoors into this paradigm and presented two types of data-poisoning-based strategies: Distributed Backdoor Attack (DBA) and Centralized Backdoor Attack (CBA) [20, 21].
  1. 1.
@@
Liu et al. extended the backdoor to node classification scenarios via a stated local loss function [22]. Additionally, they discussed various factors influencing the backdoor performance in federated graph learning (e.g., data distribution, adversary client number, trigger size, location, poisoning rate, etc.).
@@
To implant backdoor into the GCL framework, Zhang et al. first looked into the viability and proposed three methods according to the attack scenarios: (i) data-poisoning attack; (ii) set-crafting attack; (iii) natural backdoor attack [23].
  1. 1.
    Data-poisoning Attack. This attack confines adversaries to control the training data collection process of GCL.
    One core concept of GCL involves training to maximize the similarity between embeddings of the same graph node across different mapped views. Hence, by selecting a subset of nodes from the target class in the graph and connecting them to nodes with triggers, certain connections may disappear in some views after mapping through the encoder (nodes won’t aggregate trigger node information), while others will remain (nodes will aggregate trigger node information). Consequently, when calculating the loss, the encoder training aims will maximize the similarity between embeddings with trigger information and those without, hence achieving the effect of a backdoor.
  2. 2.
@@
Research [24] initially investigated the GNN backdoor in integrated circuits, utilizing subcircuits as triggers to infect the circuit dataset used for training. The embedded subcircuits do not affect the performance of the original circuit and can pass functional convergence tests. The backdoored GNN physical system classifies the circuit with triggers as harmful. Experimental results based on Hardware Trojan Detection Systems and Intellectual Property Piracy Detection Systems demonstrate that the proposed backdoor scheme exhibits excellent performance.
III-B2 Effectiveness-Improving Research.
@@
Xu et al. proposed an Explainability-based Backdoor Attack, exploiting node features with either the strongest or weakest characteristics to design subgraph triggers for poisoning the training data and embedding backdoors [25]. Executing attacks using data with prominent features makes the GNN model efficiently learn the corresponding trigger features, and facilitates more effective backdoor implantation. Conversely, launching backdoor attacks using data with weaker features minimally affects the feature learning process for normal samples, rendering the attack more covert and less prone to detection.
@@
Zheng et al. researched the distribution patterns in statistical terms between trigger subgraphs and normal subgraphs [26]. Their findings indicated that utilizing subgraphs divergent from the distribution of normal samples as triggers resulted in a notable enhancement in attack performance. Moreover, based on this phenomenon, they proposed a motif-based backdoor poisoning technique. This method employs available data to select significantly skewed subgraph distributions as motif triggers and applies node importance and subscore to determine the optimal trigger injecting position with the highest values. Then, the adversary can poison the training data with the designed trigger and backdoor the associated GNN models via learning.
Chen et al. introduced a general GNN backdoor tailored for node-level tasks [27]. Employing a trigger generator, the attack generates constrained node features specifically designed for poisoned targets, with both the GNN model and the generator being jointly trained via the poisonous dataset, which resultantly creates a backdoored GNN. Furthermore, to reduce the impact of other nodes, it selectively eliminates critical edges of targets, identified through explainability analysis. Additionally, the trigger feature dimensions crucial for targets are prioritized.
Research [28] proposed an explanatory subgraph-based poisoning attack. The subgraph incorporates graph structure and node feature information relevant to the target label, and this strengthens the association between the subgraph and the target while minimally altering the original data’s structure, and preserving its statistical properties. Also, the adversary will modify the explanation information of the graph, which makes the change of the graph more covert and reasonable.
@@
Dai et al. developed imperceptible trojan backdoor attacks under limited attack budgets [29]. Utilizing a trigger generator, the generation of infective samples ensures a sufficiently high similarity with poisoned samples from the target class data. Additionally, to make the model maintain the training accuracy on normal samples, a bi-optimization loss is designed to train both trigger generator and GNN. Furthermore, for the selection of the target for the attack, the method leverages the obtained node embeddings for clustering and chooses the most representative nodes as targets to assure backdoor efficiency.
To achieve attack concealment in graphs, study [30] developed a clean-graph backdoor attack. This attack refrains from modifying the graph’s features or structure and does not introduce malicious nodes and edges prior to training. Instead, adversaries solely alter the labels of victim nodes possessing adversary-desired distinctive features. By doing so, they ensure that the GNN training process leads the model to produce target outcomes when presented with input data exhibiting these features.
@@
Xu et al. initially delved into the feasibility of implementing clean-label backdoor attacks in GNNs [31]. It samples victim targets within the training dataset pertaining to the target class, inserting triggers within them, and then training the GNN model using both poisoned and regular data. Since the infected data belongs to the target class, it suffices to embed triggers in these samples to enable the GNN to learn trigger features without label revision. During the testing phase, the GNN will classify trigger-contained inputs as belonging to the target class. To enhance this method, Xing et al. select robust features of nodes with larger degrees as triggers [32]. These features are typically strongly associated with their respective labels. Utilizing them as triggers facilitates the model in establishing a strong correlation between the trigger and the target label.
Yang et al. evaluated the potential of clean-label attacks for node classification [33]. To realize the backdoor effect through poisoning without label modification, they embedded triggers within the target node data and introduced perturbations therein. This action causes the decision boundary of the target category to lean toward infected data during training, and leads the model to predict attacked data as the target class during the testing phase. Simultaneously, to enhance the normal performance robustness of GNN, they also proposed a contrastive poisoning strategy to slightly poison normal samples.
@@
The transferable GNN backdoor attack is proposed in [34], utilizing available real-world data to create triggers for poisoning and implanting a backdoor in a shadow GNN model via training. The mode of subgraph triggers is optimized during training using loss function derivatives to maximize the attack success rate, and it could be applied in poisoning various downstream task GNN models (i.e., transferability). Additionally, due to the harsh computing of optimizing graph structure derivatives, a solution strategy for graph structure perturbations was also presented in the study.
@@
Inspired by frequency domain graph theory and frequency-based backdoor attack methods, Zhao et al. proposed an effective spectral graph backdoor attack method [35]. It injects subtle trigger signals into the crucial frequency bands of important node features to poison training samples and implant backdoors within victim GNN. For the trigger, specifically, it first extracts the ego graph constructed from the attributes of the neighboring nodes of the target node and performs spectrum decomposition on this ego graph. Then, the adversaries inject a small trigger signal into the spectrum graph after the spectrum decomposition. The injection is done at the important frequency bands of the important features, to guarantee an efficient backdoor effect.
Graph Node Feature Backdoor. A novel backdoor approach is proposed based on graph node features, aiming to infect nodes and optimize the smoothness of graph features to generate an infected graph for backdooring GNNs [36]. Concretely, the attack employs a mask to conceal node feature data and derive features. Following that, a graph feature optimization function is utilized to modify the graph structure, resulting in the smoothest infected graph containing the trigger. The features of the graph are smoother, minimizing the impact on feature learning from the original data. The backdoor keeps better preservation of the model’s accuracy on normal samples while ensuring the robustness of the backdoor performance.
Graph Link Backdoor. Link prediction is an essential task in GNN, primarily targeting at predicting links (edges or connections) that are not yet present but may appear in the future within a graph. Zheng et al. initially analyzed the feasibility of implementing backdoors in such tasks, exploiting poisoning to backdoor GNN [37]. It first designates a specified subgraph structure as the trigger, then optimizes it using gradient information from the prediction model, and finally poisons the training data with the optimized trigger to backdoor the victim GNN. Dai et al. further improved the trigger by leveraging nodes as the trigger. During training, edges between nodes connected to this trigger are misclassified into an adversary-specified target [38].
@@
Study [39] reviewed the possibility of implementing such attacks within graph learning framework. First, explainability techniques are applied to determine multi-trigger locations and identify poisoned graph nodes according to these locations. Next, a one-hot encoding mechanism is used to encode the poisoned nodes and multiple target labels. Finally, one carefully designed loss function, coupled with a two-tier-optimization mechanism, is leveraged to train the multi-trigger generation model and victim GNN, resulting in the GNN backdoored that responds to multi-triggers after learning.
Xu et al. implemented multi-target backdoors through multi-target data-poisoning [40]. The attack primarily centers on two backdoor strategies: One-to-N, in which distinct numerical values of several feature dimensions regulate various target outcomes, and N-to-One, whereby when all the triggers are satisfied, the backdoor could be activated.
TABLE II: Summary of Graph Neural Backdoor Defense Strategies
@@
BloGBaD (2024) [41] $\bullet$ - - $\bullet$ - $\bullet$ $\diamond$ Backdoor detection $\diamond$ Sample filtration $\diamond$ Backdoor mitigation Employ sample distributions to identify infected graphs
Explainability Defense (2022) [42] $\bullet$ - - $\bullet$ $\bullet$ - $\diamond$ Backdoor detection $\diamond$ Sample filtration Leverage explainability tools to detect poisonous graphs
Securing GNN (2024) [43] $\bullet$ - - $\bullet$ $\bullet$ - $\diamond$ Backdoor detection Use graph features to detect sample anomalies
III-C Defense Methodologies
In response to the threat of GNN backdoor attacks, researchers have begun focusing on developing effective defense strategies. The comprehensive summary of GNN backdoor defense studies is displayed in Tab. II.